Return to question

Item2

Item 2 Item Skipped

A rotary engine powers a vertical takeoff and landing (VTOL) personal aircraft known as the Moller Skycar M400. It is a flying car known as a personal air vehicle (PAV), and it is expected to make its first untethered flight in 2020. The PAV has been under development for 25 years at a total cost of $90 million. Assuming the $90 million was spent in an equal amount each year, determine the future worth at the end of the 25-year period at an interest rate of 9% per year.

The future worth is $
Answer #1

F=A(F/A,i,n)

90 million will be spent in an equal amount each year so, A= 90000000/25 = 3600000

i=0.09

n=25

F = 3600000[(1+0.09)^25-1/0.09]

3600000[7.623/0.09]

3600000*84.7

304920000